Principles of Advanced Taxation Planning

1. What payments would a company make this FBT year if it had a $5,200 FBT liability for the current year and they had a $5,000 FBT liability last FBT year?

2. Bonny Tyler receives the following in the current year:
Salary and wages $53,000
Fully franked dividends $4,500
Unfranked dividends $500
How much tax and Medicare will Bonny have to pay given she has private health insurance?

3. If a business buys a product for $110 including GST and sells it for $220 including GST, how much GST will the business end up paying?

4. How would financial advisers use information relating to GST to assist their clients?
Answer #1

1 As FBT payment is made in Advance so During Current period on $5200 liability arises.

2. She Needs to Pay only on Unfranked Dividend as No tax has been deducted by the company on the same so taxable in the hands of Sharrholder.

Salary and wages below Taxable Slab so not required to pay tax as she has Health Insurance also.

Franked Dividend are not taxable as company Deducted tax on the same.

3. Suppose GSt Rate is 10%

NET GST Liablity=20-10=10

Working Note:1

Purchase Price (Including GST)=110

BAsic Purchase Price=110/(100+10)*100=100

Input GSt on Purchase=110-100=10

Working Note:2

Sale PRice(Including GST)=220

BAsic Sale Price=220/(100+10)*100=200

GSt Liablity on Sale=220-200=20

4. GST help Financial adviser to Evaluate theri true & Actual GST liablity, GST Input & on the basis of  Sale of Goods & services , Financial Adviser helps the Stakeholders to calculate their Estimated Tax Liablity And Advance TAx .